OpenStudy (anonymous):

Differentiate function y=(square root of x)(x-1)

OpenStudy (anonymous):

have you thought of expanding?

OpenStudy (anonymous):

well I got x as an answer

OpenStudy (anonymous):

but maybe i did something wrong, should of turned x^1/2 and multiplied or use the rule?

OpenStudy (anonymous):

i dont believe that is right

OpenStudy (anonymous):

try expanding it or use the product rule, basically amount to the same thing though

OpenStudy (anonymous):

im still stuck

OpenStudy (anonymous):

\[\sqrt{x}(x-1)=x^{\frac{3}{2}}-x^{\frac{1}{2}}\] and now power rule

OpenStudy (mimi_x3):